# The Road to Trust: Building Enclaves within Confidential VMs

### v2

@@ 6.3. Real-world evaluations
Figure 7. Real world application benchmarks.
We conducted evaluations on 4 real-world applications, including hash join (memory- and computation-intensive tasks), SQLite (I/O- and memory-intensive tasks) and TLS server (I/O-intensive tasks) which ran directly on NestedSGX, and Redis (I/O- and memory-intensive tasks) which ran with Occlum on NestedSGX.
@@
Hash join. Hash join is used to implement the “equivalent-join” operation in modern databases. It involves creating a hash table from rows of the first table and then probes it with rows of the second table. We used the open-source implementation of the algorithm in SGXGauge (Kumar et al., 2022; sgx, 2023b). We varied the size of the first table (from 500K to 1M records) and fixed the second table (100K), effectively varying the memory and computation intensive nature of the workload. As shown in Figure. 7(a), the overhead of NestedSGX is negligible.
@@
SQLite. We utilized the publicly available version of SQLite (v3.19.3) operating on Intel SGX (sql, 2021). To evaluate the memory performance of NestedSGX, we configured the database to function in-memory mode and incorporated the client within the enclave. We conducted performance evaluations across various record sizes, measuring the time required for 100,000 database operations. Specifically, we employed three representative workloads within the Yahoo! Cloud Serving Benchmark (YCSB) suite (ycs, 2021): workload $A$ (update heavy with 50% reads and 50% updates), $D$ (read latest, i.e., delete old ones, insert new ones and read mostly the new ones), and $F$ (short range scan) for the evaluation. As shown in Figure 7(b), NestedSGX introduces about 0.77% overhead on average over the baseline.
@@
TLS server. We utilized the open-source SGX-OpenSSL project and a sample implementation of a TLS server in an enclave (sgx, 2023a) for our evaluation. The TLS client operated within the native CVM and sent requests to the server over the local loop-back. To measure the average latency, we employed the tls-perf benchmarking tool (tls, 2023) and set the number of requesting threads to 1, 2, 5, 10, 20, 30, 40, 50 respectively. As illustrated in Figure 7(c), NestedSGX incurs approximately a 9.75% overhead over the baseline.
@@
Redis. We ran a Redis (v6.0.9) database server with Occlum on NestedSGX. The database was configured as an in-memory setup, and we applied the YCSB workload A for our evaluation. We first loaded 5,000 records (in total 5 MB data), and subsequently executed 10,000 operations from 20 clients over the local loopback. We gradually increased the request frequency and measured the corresponding latency at different throughput levels. As depicted in Figure 7(d), NestedSGX achieved an approximate throughput of 84.32% (15.68% overhead) compared to the baseline. The latency was found to be 6.4% under low load conditions and increased to approximately 55% as the system approached its maximum throughput.
7. Discussions
@@
Supporting other VM-based TEEs. Recently, ARM CCA introduces the support of different planes within a CVM, in which each plane is essentially a separated VM but they all share the same guest physical address space. Plane 0 is more privileged and can run a paravisor which manages the switches between other planes and can restrict the memory accessible by other planes (Arm, 2024). According to ARM’s roadmap, the paravisor is designed to support secure services such as vTPM emulation. Similarly, we can re-purpose the paravisor to support running userspace enclaves. Similarly, TDX supports memory sharing among CVMs with the assistance of the TDX module. Therefore, it is promising that NestedSGX can be extended to other CVM platforms. However, it is worth mentioning that the TDX module and CCA’s Realm Management Monitor (RMM) are critical for the memory isolation and secure management of CVMs, and the customization of these components may not be allowed by chip companies. In fact, only the TDX module signed by Intel can run in the Secure Arbitration Mode (SEAM).
@@
Supporting other TEE abstractions. NestedSGX adopts the SGX model as it is arguably the most prevalent TEE with process abstraction. The NestedSGX platform can be enhanced to accommodate other TEE abstractions, such as ARM TrustZone. Specifically, VMPL0 could run the trusted OS (e.g., OP-TEE) and trusted applications (TAs), providing secure services to the guest OS running at VMPL1, which is analogous to the normal world.
The SMC instruction is used to handle transitions between the secure world and normal world. In TrustZone, the SMC handler is within the trusted firmware. In our system, the SMC instruction needs to be replaced with a VMPL switching request, which transfers the execution to the trusted OS. This allows porting existing applications built on ARM TrustZone to SEV-SNP CVMs.
8. Related Works
@@
Integrity Measurement Architecture (IMA). Integrity plays a vital role in ensuring system security by guaranteeing the exclusive loading of authentic software onto a machine. Measured boot, utilizing the Trusted Platform Module (TPM), securely captures measurements of all boot software, culminating in the loading and execution of the OS kernel. The Linux Integrity Measurement Architecture (IMA) expands upon the concept of measured boot by comprehensively recording all software executions and file accesses within the OS, securely storing them in the TPM (Sailer et al., 2004). The Container-IMA (C-IMA) extends IMA to containers, enabling the measurement of container images and the runtime integrity measurement of container processes (Luo et al., 2019).
@@
In lieu of TPM, TZ-IMA extends the storage of measurements in TrustZone (Song et al., 2022). Recently, Intel introduced a solution of IMA in confidential cloud environments, with a focus on Intel TDX, enabling IMA of programs and applications running in Intel trusted domain (TD) during runtime (tdi, 2023).
@@
However, IMA in general assumes the trust of the boot components, which includes the BIOS, bootloader and guest OS, leading to potentially large TCB and expanded attack surface. For example, it has been shown that IMA can be bypassed with a malicious block device (Bohling et al., 2020). In contrast, the core root of trust for measurement of NestedSGX (i.e., the security monitor) operates at VMPL0 and we do not place trust on the guest OS. Moreover, the security monitor not only measures the integrity of both enclave code and data but also ensures that the enclave runs in an isolated environment from the guest OS, which is beyond the design goal of IMA.
@@
Establishing the enclave abstraction within TEEs. With the emergence of Intel SGX, endeavors have been undertaken to enhance the interoperability between Intel SGX and other TEE platforms. Notably, Komodo (Ferraiuolo et al., 2017), Sanctuary (Brasser et al., 2019), and SecTEE (Zhao et al., 2019) offer support for enclave abstractions on ARM TrustZone. HyperEnclave (Jia et al., 2022) re-introduces Intel SGX on standard AMD hardware through the use of a small and trusted hypervisor. In contrast to these works, NestedSGX sets itself apart by adopting a distinctive model. While the guest OS remains shielded from the untrusted hypervisor, it is not entirely trusted. Consequently, NestedSGX enables enhanced protection for enclaves within the feature-rich guest OS. Most related to our work are vSGX (Zhao et al., 2022) and Veil (Ahmad et al., 2024). A more detailed comparison with them is provided in Sec. 3.1.
@@
Security and performance enhancement using the VMPL feature. Hecate (Ge et al., 2022) utilizes VMPL0 as an L1 hypervisor, operating within the CVM on top of the untrusted L0-hypervisor (host VMM). Hecate enables the migration of on-premises workloads by maintaining compatibility with unmodified VM images and implementing security policies, such as network firewalls. Honeycomb (Mai et al., 2023) executes a validator within VMPL0, safeguarded from other software. The validator analyzes the binary code of a GPU kernel to validate that each memory instruction within the GPU kernel can only access specific regions of the virtual address space, employing static analysis techniques. SVSM-vTPM (Narayanan et al., 2023) employs VMPL to isolate the virtual TPM (vTPM) from the guest OS, ensuring the integrity of vTPM’s functionalities and the security of the root key for remote attestation. However, all these solutions do not provide isolation for user programs from the untrusted guest OS.
9. Conclusions
@@
Although confidential computing aims to protect the confidentiality and integrity of code and data within TEEs, it remains challenging to control code within CVMs since the feature-rich guest OS is free to load any code. This paper presents NestedSGX, a secure and efficient solution for maintaining control over the integrity of code and data in TEEs. NestedSGX creates hardware enclaves inside guest VMs with VMPL, which ensures remote attestation of trusted and measured code. NestedSGX is compatible with Intel SGX and easily integrated into existing systems. Our evaluation shows NestedSGX incurs small overhead in most real-world applications.

### v3

@@ VI-C Real-world evaluations
Figure 7: Real world application benchmarks.
@@
Hash join. Hash join is used to implement the “equivalent-join” operation in modern databases. It involves creating a hash table from rows of the first table and then probes it with rows of the second table. We used the open-source implementation of the algorithm in SGXGauge [34, 8]. We varied the size of the first table (from 500K to 1M records) and fixed the second table (100K), effectively varying the memory and computation intensive nature of the workload. As shown in Fig. 7(a), the overhead of NestedSGX is negligible.
@@
SQLite. We utilized the publicly available version of SQLite (v3.19.3) operating on Intel SGX [3]. To evaluate the memory performance of NestedSGX, we configured the database to function in-memory mode and incorporated the client within the enclave. We conducted performance evaluations across various record sizes, measuring the time required for 100,000 database operations. Specifically, we employed three representative workloads within the Yahoo! Cloud Serving Benchmark (YCSB) suite [2]: workload $A$ (update heavy with 50% reads and 50% updates), $D$ (read latest, i.e., delete old ones, insert new ones and read mostly the new ones), and $F$ (short range scan) for the evaluation. As shown in Fig. 7(b), NestedSGX introduces about 0.77% overhead on average over the baseline.
@@
TLS server. We utilized the open-source SGX-OpenSSL project and a sample implementation of a TLS server in an enclave [6] for our evaluation. The TLS client operated within the native CVM and sent requests to the server over the local loop-back. To measure the average latency, we employed the tls-perf benchmarking tool [9] and set the number of requesting threads to 1, 2, 5, 10, 20, 30, 40, 50 respectively. As illustrated in Fig. 7(c), NestedSGX incurs approximately a 9.75% overhead over the baseline.
@@
Redis. We ran a Redis (v6.0.9) database server with Occlum on NestedSGX. The database was configured as an in-memory setup, and we applied the YCSB workload A for our evaluation. We first loaded 5,000 records (in total 5 MB data), and subsequently executed 10,000 operations from 20 clients over the local loopback. We gradually increased the request frequency and measured the corresponding latency at different throughput levels. As shown in Fig. 7(d), NestedSGX achieved a throughput of 84.32% (15.68% overhead) compared to the baseline. The latency overhead was found to be 6.4% under low load conditions and increased to approximately 55% as the system approached its maximum throughput.
VII Limitations & Future Works
Limitations. Since our implementation is still in the prototype stage, NestedSGX has some limitations to fully support the SGX model. Firstly, the security monitor manages the enclave’s page tables, which prevents the guest OS from manipulating the enclave’s page tables for potential address mapping attacks. Since the App and enclave have separate page tables, the enclave cannot access the App’s memory, as their memory views may differ if the App’s page table is updated without synchronizing with the enclave’s page table. Therefore, features such as user_check and memory sharing are not currently supported by NestedSGX. Secondly, the NestedSGX-driver does not currently support swapping the enclave’s memory to disks. Thirdly, the guest OS lacks the capability to inject an interrupt into VMPL0. This limitation imposes constraints on the OS’s scheduling of enclave threads. Finally, compatibility is currently provided on top of SGX SDK and Occlum. Thus, applications that are purely written in the SGX instruction set without any SDK or LibOS are not yet supported. We do not think that any of these limitations are inherent to NestedSGX, e.g., memory sharing could be supported if updates to the App’s page table were synchronized with the enclave’s page table via VMPL switches.
@@ VII Limitations & Future Works
Supporting other VM-based TEEs. Recently, ARM CCA introduces the support of different planes within a CVM, in which each plane is essentially a separated VM but they all share the same guest physical address space. Plane 0 is more privileged and can run a paravisor which manages the switches between other planes and can restrict the memory accessible by other planes [18]. According to ARM’s roadmap, the paravisor is designed to support secure services such as vTPM emulation. Similarly, we can re-purpose the paravisor to support running userspace enclaves. TDX supports a similar feature named TD partitioning. Therefore, it is promising that NestedSGX can be extended to other CVM platforms. However, it is worth mentioning that the TDX module and CCA’s Realm Management Monitor (RMM) are critical for the memory isolation and secure management of CVMs, and the customization of these components may not be allowed by chip companies. In fact, only the TDX module signed by Intel can run in the Secure Arbitration Mode (SEAM).
@@
Supporting other TEE abstractions. NestedSGX adopts the SGX model as it is arguably the most prevalent TEE with process abstraction. The NestedSGX platform can be enhanced to accommodate other TEE abstractions, such as ARM TrustZone. Specifically, VMPL0 could run the trusted OS (e.g., OP-TEE) and trusted applications (TAs), providing secure services to the guest OS running at VMPL1, which is analogous to the normal world. The SMC instruction is used to handle transitions between the secure world and normal world. In TrustZone, the SMC handler is within the trusted firmware. In our system, the SMC instruction needs to be replaced with a VMPL switching request, which transfers the execution to the trusted OS. This allows porting existing applications built on ARM TrustZone to SEV-SNP CVMs.
VIII Related Works
@@
Integrity Measurement Architecture (IMA). Integrity plays a vital role in ensuring system security by guaranteeing the exclusive loading of authentic software onto a machine. Measured boot, utilizing the Trusted Platform Module (TPM), securely captures measurements of all boot software, culminating in the loading and execution of the OS kernel. The Linux Integrity Measurement Architecture (IMA) expands upon the concept of measured boot by comprehensively recording all software executions and file accesses within the OS, securely storing them in the TPM [43]. The Container-IMA (C-IMA) extends IMA to containers, enabling the measurement of container images and the runtime integrity measurement of container processes [36].
@@
In lieu of TPM, TZ-IMA extends the storage of measurements in TrustZone [49]. Recently, Intel introduced a solution of IMA in confidential cloud environments, with a focus on Intel TDX, enabling IMA of programs and applications running in Intel trusted domain (TD) during runtime [7].
@@
However, IMA in general assumes the trust of the boot components, which includes the BIOS, bootloader and guest OS, leading to potentially large TCB and expanded attack surface. For example, it has been shown that IMA can be bypassed with a malicious block device [20]. In contrast, the core root of trust for measurement of NestedSGX (i.e., the security monitor) operates at VMPL0 and we do not place trust on the guest OS. Moreover, the security monitor not only measures the integrity of both enclave code and data but also ensures that the enclave runs in an isolated environment from the guest OS, which is beyond the design goal of IMA.
@@
Establishing the enclave abstraction within TEEs. With the emergence of Intel SGX, endeavors have been undertaken to enhance the interoperability between Intel SGX and other TEE platforms. Notably, Komodo [24], Sanctuary [21], and SecTEE [61] offer support for enclave abstractions on ARM TrustZone. HyperEnclave [31] re-introduces Intel SGX on standard AMD hardware through the use of a small and trusted hypervisor. In contrast to these works, NestedSGX sets itself apart by adopting a distinctive model. While the guest OS remains shielded from the untrusted hypervisor, it is not entirely trusted. Consequently, NestedSGX enables enhanced protection for enclaves within the feature-rich guest OS. Most related to our work are vSGX [62] and Veil [13]. A more detailed comparison with them is provided in Sec. III-A.
@@
Security and performance enhancement using the VMPL feature. Hecate [25] utilizes VMPL0 as an L1 hypervisor, operating within the CVM on top of the untrusted L0-hypervisor (host VMM). Hecate enables the migration of on-premises workloads by maintaining compatibility with unmodified VM images and implementing security policies, such as network firewalls. Honeycomb [37] executes a validator within VMPL0, safeguarded from other software. The validator analyzes the binary code of a GPU kernel to validate that each memory instruction within the GPU kernel can only access specific memory regions, employing static analysis techniques. SVSM-vTPM [40] employs VMPL to isolate the virtual TPM (vTPM) from the guest OS, ensuring the integrity of vTPM’s functionalities and the security of the root key for remote attestation. However, all these solutions do not provide isolation for user programs from the untrusted guest OS.
IX Conclusions
@@
Although confidential computing aims to protect the confidentiality and integrity of code and data within TEEs, it remains challenging to control code within CVMs since the feature-rich guest OS is free to load any code. This paper presents NestedSGX, a secure and efficient solution for maintaining control over the integrity of code and data in TEEs. NestedSGX creates hardware enclaves inside CVMs with VMPL, which ensures remote attestation of trusted and measured code. NestedSGX is compatible with Intel SGX and easily integrated into existing systems. NestedSGX incurs a small overhead in most real-world applications.
Acknowledgments
We would like to express our sincere gratitude to the anonymous reviewers and our shepherd. The project was supported in part by the National Key R&D Program of China (Grant No. 2020YFB1805402), the National Natural Science Foundation of China (Grant No.62272452) and the research grant from Ant Group. Corresponding authors: Shoumeng Yan (shoumeng.ysm@antgroup.com) and Rui Hou (hourui@iie.ac.cn).